An 'α' particle has an initial velocity of (12 km/s)ˆj+ (15 km/s)ˆkand aconstant acceleration of (2×108m/s2)ˆi in a region where there are uniform electricand magnetic fields. If~B= (4×10−4,0,0) T, what is~E?


1
Expert's answer
2021-03-09T10:58:38-0500
"\bold{F =ma= qE + qv × B}\\\bold{v × B}=(12000,15000,0)×(4\cdot10^{-4},0,0)\\=(0,0,-6)"

"(6.64\cdot10^{-27})(2\cdot10^8,0,0)= (3.22\cdot10^{-19})E \\+ (3.22\cdot10^{-19})(0,0,-6)"

"\bold{E}=(4.1,0,6)\frac{V}{m}"